1. True or False: When the overall average level of prices of goods and services falls, the inflation rate falls.

2. Study Questions and Problems #2

The following table shows the value of the market basket of goods in years 2015, 2016, and 2017.

Year

Value of Market Basket

(Dollars)

2015 205
2016 240
2017 255

Suppose that the base year is 2015.

The value of the CPI in 2016 is approximately .

Continue to assume that the base year is 2015.

The value of the CPI in 2017 is approximately .

The rate of inflation between 2016 and 2017 is approximately %.

3.

Suppose you earned $130,000 in 2014.

If the CPI is 125 for 2014, the value of your real income is $
Answer #1

1. True

2. CPI in 2016 = 240*100/205= 117.073

CPI in 2017= 255*100/205= 124.39

Inflation rate = (124.39-117.073)*100/117.073= 6.25

3. Value of real income = 130000*100/125= 104000